# A Galactic Transformation—Understanding the SMC’s Structural and Kinematic Disequilibrium

### v1

@@ 4.1 Impact of LMC Ram Pressure on the SMC’s Gas
Note that exploring LMC’s ram pressure in detail is not within the scope of this paper. Further, the SPH nature of the B12 simulation limits the ability to fully resolve the hydrodynamic forces like ram pressure (Agertz2007; Sijacki2012; Hayward2014), and more advanced mesh-based simulations with high spatial resolution and a truly multi-phase ISM are needed, in order to properly resolve the physical processes at fluid interfaces (like the Kelvin-Helmholtz instability). However, here we present some order of magnitude analytical calculations to help understand the role of ram pressure.
We use the analytical formulation of Gunn1972 to estimate the relative strength of the LMC’s ram pressure and the SMC’s own gravitational restoring thrust on the SMC’s gas. The GG72 formulation is routinely used (e.g. Roediger2005; Roediger2006; Tonnesen2009; Salem2015) to obtain order of magnitude estimates of the ram pressure experienced by satellite galaxies as they infall into a circumgalactic medium (CGM) or an intra-cluster medium (ICM). We apply the GG72 formulation to the SMC’s gas as it passes through the LMC’s ISM during a recent collision, as described below.
@@
Following Salem2015, the SMC’s gravitational restoring thrust (restoring force per unit area; $P_{SMC}$), at a radial location $R$ inside its disk, is computed as:
@@
Figure 13 shows $P_{\rm ram}$ and $P_{\rm SMC}(R)$. The ram pressure is more than an order of magnitude larger than the restoring thrust at all radii. In the above arguments, we have not accounted for the ram pressure on the SMC’s gas from the MW’s CGM. Further, Ultra-Violet absorption spectroscopic surveys suggest the presence of a hot (at least $\sim 10^{4}$ K) gas component associated with the LMC as well (Krishnarao2022). However, ram pressure from a host galaxy’s CGM on a satellite ranges between $\mathcal{O}(10^{2})-\mathcal{O}(10^{3})$ M kpc Myr-2 (Salem2015; Emerick2016), whereas the ram pressure from the LMC’s gas disk on the SMC is much larger ($\mathcal{O}(10^{5})$ M kpc Myr-2). Only extreme extra-galactic environments can lead to such large ram pressures, like the central regions of massive galaxy clusters ($10^{15}$ M) where galaxies move at speeds exceeding 1000 km s-1 (Xie2025).
@@
This calculation highlights a physical effect that is not generally considered in studies of satellite galaxy evolution. Theoretical studies find it challenging to remove enough gas from the inner regions of satellites in order to explain their quenching timescales of $\sim 2$ Gyr (e.g. Emerick2016). Here, we illustrate that collisions between low mass galaxies (e.g. pre-processing in low mass groups) can help provide an extra kick to the gas, which aids in subsequent gas removal through other environmental processes (e.g. ram pressure from the host CGM, host tidal field). Hence, such ram pressure kicks during galaxy collisions are expected to accelerate dIrr to dE/dSph transition. The strength of ram pressure is equivalent to an ICM environment, where dEs are known to exist. Further, these kicks are impulsive, and hence they are particularly significant for low mass galaxies with a shallow potential depth. Additionally, such ram pressure kicks from the primary galaxy’s ISM during close satellite interactions need to be taken into account for interpreting the existence of dEs around M31 (e.g. NGC 147, NGC 185, NGC 205) and the evolution of Sagittarius dSph as it orbits around the MW.
4.2 The SMC’s Gas LoS Distribution
@@
Figure 14: The distribution of gas as a function of distance along the longest axis of the stellar distribution in the Model 2 simulated SMC. A distance of 0 kpc on the x-axis corresponds to the SMC’s stellar density center, and positive values are in the direction of the tidal tail. Gas particles are selected to reside within 4 kpc cylinder about the longest axis. Three epochs are shown - 100 Myr (left panel), 150 Myr (middle panel) and 200 Myr (right panel) post SMC-LMC collision. At all times, the gas distribution peaks at the stellar density center. 100 Myr post-collision, the gas distribution has a prominent secondary peak at 3-4 kpc. This secondary peak is due to a gas tidal tail that develops from the SMC-LMC collision. With time, the amplitude of the secondary peak decreases and the gas distribution in the tidal tail becomes more elongated. The gas present along the SMC’s tidal tail needs to be taken into account when interpreting the observed SMC’s HI LoS distance bi-modality (Murray2024).
@@
As mentioned in section 1, several observational studies indicate that the SMC’s HI shows two prominent velocity components along the LoS. These components have similar brightness temperatures, indicating they have similar HI mass. Murray2024 selected a sample of young stars (age << 10 Myr) from the Gaia and APOGEE surveys that have similar LoS velocities as the HI components. After accounting for dust extinction, they find that the stars associated with the HI components are separated by 5 - 10 kpc along the LoS. Hence, the SMC’s HI LoS velocity bimodality is likely due to a bimodality in the gas distribution along the LoS. In this section, using the B12 Model 2 simulation, we attempt to provide an interpretation of the complex LoS structure of the SMC’s ISM.
@@
We showed in section 3.1 that the SMC’s stellar tidal tail born out of the SMC-LMC collision can explain the observed SMC’s stellar LoS depth. The gas will also form a tidal tail similar to the stars. We investigate whether the elongated structure of the SMC’s gas tidal tail can help us interpret the observed HI distance bimodality.
@@
The Model 2 simulated SMC’s gas distribution is centered at the stellar density center. The coordinate axes are aligned with the principal axes of the stellar distribution (see section 3.1). As such, the longest stellar axis is used as a proxy for the LoS.
@@
Hence, the gas distribution along the SMC’s tidal tail needs to be taken into account when interpreting the LoS structure of the observed SMC’s HI. Such a tidal gas tail can result in a bimodal gas distribution along the LoS, where the primary gas component corresponds to the SMC’s main body, and the secondary component corresponds to the extended gas distribution along the tail.
@@
4.3 Implications for Dynamical Mass Measurements of the SMC
@@
Precise measurement of the SMC’s DM content is important for several purposes. These include constraining the SMC-LMC-MW orbit (Kallivayalil2013; Zivick2018; Patel2020) and understanding the effect of the SMC on the LMC’s disk structure (e.g. Pardy2016; Arranz2025; Rathore2025b), internal kinematics (Choi2022), star formation history (Massana2022; Harris2009), and dark matter halo (H. Foote et al. in prep). Further, precise measurement of the SMC’s DM mass is needed to accurately constrain the low mass end of the cosmological stellar mass - halo mass relation (e.g. Wechsler2002; Behroozi2019; Bowden2023) and to ultimately use the SMC-LMC-MW system as a testbed for DM physics (Foote2023).
@@
Equilibrium is typically assumed in the majority of existing measurements of the SMC’s dynamical mass. For example, studies have assumed that the SMC has a stable rotation curve (Stanimirovic1999; Stanimirovic2004; Bekki2009; DiTeodoro2019) or that the Virial mass estimator is valid (Dopita1985; Harris2006).
If the SMC is assumed to be a rotating disk, rotation curves can be fit to the velocity data to derive a mass profile. For example, DiTeodoro2019 applied this method to derive a total mass of total mass of $\sim 2.4\times 10^{9}$ M within 4 kpc of the SMC’s HI kinematic center. If the SMC is assumed to be in virial equilibrium, virial mass estimators can be applied to the stellar velocity data to derive a dynamical mass. For example, HZ06 used LoS velocities of red giant stars and determined the SMC’s total enclosed mass within $\approx 2$ kpc to be $1.4-1.9\times 10^{9}$ M, and within $\approx 3$ kpc to be $2.7-5.1\times 10^{9}$ M.
@@
Virial mass estimators have been shown to reliably recover the mass profile for galaxies facing significant external perturbations and interactions (e.g. Lynden-Bell1967; Silva2017; Hamilton2024). This is because these galaxies achieve Virial equilibrium within a few dynamical timescales. However, the SMC-LMC collision occurred $\sim$100 Myr ago, which is less than a dynamical time, even in the inner regions of the SMC (Rathore2025b). This fact calls to question the validity of using Virial mass estimators in the SMC.
@@
However, post-collision, the ratio varies by a factor of 3. The uncertainty will be higher when applied to observational data, as $f$ cannot be calibrated. For instance, if $f$ itself is uncertain to a factor of 2, the uncertainty in the observed SMC’s Virial mass estimates can be up to a factor of 6. This level of uncertainty is sufficient to change the response of the LMC disk to the SMC’s collision (e.g., the LMC’s bar tilt Rathore2025b) and the SMC’s orbital history (Arranz2024). Fundamentally, this level of uncertainty is not sufficient to constrain the inner DM density profile of the SMC, which is a critical test of DM physics (Rathore2025b).
@@
To precisely measure the SMC’s DM content, we need approaches that do not rely on equilibrium assumptions. A way forward is to use the SMC’s observable perturbations on the LMC’s disk during the SMC-LMC collision to measure the SMC’s DM content. R25 showed that morphological perturbations in the LMC, such as the LMC’s bar tilt ($5^{\circ}-15^{\circ}$) (Choi2018; Arranz2025b), are caused by the SMC’s torques during the collision. R25 estimated the SMC’s pre-collision total enclosed mass within $R_{\rm enc}=2$ kpc to be $0.8-2.4\times 10^{9}$ M by modeling the SMC’s gravitational torques on the LMC’s bar. Further, they suggest that the precision in the SMC’s mass measurement in their framework is largely limited by the observational uncertainty in the LMC bar’s tilt. Hence, it is important to get more precise observational constraints on the disequilibrium morphology and kinematics of the LMC and SMC to enable more precise determinations of the SMC’s mass, without relying on equilibrium assumptions.
4.4 Comparison with Previous Studies
Several works have presented simulations of the SMC-LMC-MW interactions (e.g. Gardiner-Noguchi1996; Connors2006; Besla2010; DB2012; Pardy2018; Cullinane2022; Lucchini2024; Arranz2024). However, the majority of the aforementioned studies have focused on either the LMC’s internal structure and kinematics, or the MS. Further, most previous studies have modeled the SMC with either static potentials or spherical halos, which limits detailed investigations of the SMC’s internal structure and kinematics as the SMC’s halo will be significantly deformed throughout its interaction with the LMC (Foote et al. in prep). To the best of our knowledge, the most recent work that has investigated the internal structure and kinematics of the SMC with simulations is DB2012. Hence, we primarily compare our analysis and results with DB12.
@@ 4.4 Comparison with Previous Studies
During the most recent SMC-LMC pericenter in DB12, two tidal features form due to the LMC’s tidal forces on the SMC. The first feature is a bridge of stars between the SMC and LMC. The second feature is a tidal tail of stars, which DB12 refer to as the “counter-bridge” (see also Connors2006). DB12 argue that the large observed LoS depth results from a viewing perspective that aligns along the “counter-bridge”. In our study, we have similarly argued for this scenario, and shown that a tidal tail can also be generated from a direct collision where the impact parameter is much smaller.
@@
With the B12 simulations, we have shown that the SMC can have a rotating old stellar disk and gas disk prior to the collision. This is consistent with the known morphology of isolated dIrr galaxies (McGaugh2021). Through a collision, strong LMC tides can transform the SMC’s rotating stellar disk to a dispersion dominated system, consistent with observations.
@@
4.5 Limitations and Future Work

### v2

@@ IV.1 Impact of LMC Ram Pressure on the SMC’s Gas
Note that exploring LMC’s ram pressure in detail is not within the scope of this paper. Further, the SPH nature of the B12 simulation limits the ability to fully resolve the hydrodynamic forces like ram pressure (Agertz et al., 2007; Sijacki et al., 2012; Hayward et al., 2014), and more advanced mesh-based simulations with high spatial resolution and a truly multi-phase ISM are needed, in order to properly resolve the physical processes at fluid interfaces (like the Kelvin-Helmholtz instability). However, here we present some order of magnitude analytical calculations to help understand the role of ram pressure.
We use the analytical formulation of Gunn and Gott (1972, hereafter GG72) to estimate the relative strength of the LMC’s ram pressure and the SMC’s own gravitational restoring thrust on the SMC’s gas. The GG72 formulation is routinely used (e.g. Roediger and Hensler, 2005; Roediger and Brüggen, 2006; Tonnesen and Bryan, 2009; Salem et al., 2015) to obtain order of magnitude estimates of the ram pressure experienced by satellite galaxies as they infall into a circumgalactic medium (CGM) or an intra-cluster medium (ICM). We apply the GG72 formulation to the SMC’s gas as it passes through the LMC’s ISM during a recent collision, as described below.
@@
Following Salem et al. (2015), the SMC’s gravitational restoring thrust (restoring force per unit area; $P_{SMC}$), at a radial location $R$ inside its disk, is computed as:
@@
Figure 13 shows $P_{\rm ram}$ and $P_{\rm SMC}(R)$. The ram pressure is more than an order of magnitude larger than the restoring thrust at all radii. In the above arguments, we have not accounted for the ram pressure on the SMC’s gas from the MW’s CGM. Further, Ultra-Violet absorption spectroscopic surveys suggest the presence of a hot (at least $\sim 10^{4}$ K) gas component associated with the LMC as well (Krishnarao et al., 2022). However, ram pressure from a host galaxy’s CGM on a satellite ranges between $\mathcal{O}(10^{2})-\mathcal{O}(10^{3})$ M kpc Myr-2 (Salem et al., 2015; Emerick et al., 2016), whereas the ram pressure from the LMC’s gas disk on the SMC is much larger ($\mathcal{O}(10^{5})$ M kpc Myr-2). Only extreme extra-galactic environments can lead to such large ram pressures, like the central regions of massive galaxy clusters ($10^{15}$ M) where galaxies move at speeds exceeding 1000 km s-1 (Xie et al., 2025).
@@
This calculation highlights a physical effect that is not generally considered in studies of satellite galaxy evolution. Theoretical studies find it challenging to remove enough gas from the inner regions of satellites in order to explain their quenching timescales of $\sim 2$ Gyr (e.g. Emerick et al., 2016). Here, we illustrate that collisions between low mass galaxies (e.g. pre-processing in low mass groups) can help provide an extra kick to the gas, which aids in subsequent gas removal through other environmental processes (e.g. ram pressure from the host CGM, host tidal field). Hence, such ram pressure kicks during galaxy collisions are expected to accelerate dIrr to dE/dSph transition. The strength of ram pressure is equivalent to an ICM environment, where dEs are known to exist. Further, these kicks are impulsive, and hence they are particularly significant for low mass galaxies with a shallow potential depth. Additionally, such ram pressure kicks from the primary galaxy’s ISM during close satellite interactions (see also Matijević et al. 2025) need to be taken into account for interpreting the existence of dEs around M31 (e.g. NGC 147, NGC 185, NGC 205) and the evolution of Sagittarius dSph as it orbits around the MW.
IV.2 The SMC’s Gas LoS Distribution
@@
Figure 14: The distribution of gas as a function of distance along the longest axis of the stellar distribution in the Model 2 simulated SMC. A distance of 0 kpc on the x-axis corresponds to the SMC’s stellar density center, and positive values are in the direction of the tidal tail. Gas particles are selected to reside within 4 kpc cylinder about the longest axis. Three epochs are shown - 100 Myr (left panel), 150 Myr (middle panel) and 200 Myr (right panel) post SMC-LMC collision. At all times, the gas distribution peaks at the stellar density center. 100 Myr post-collision, the gas distribution has a prominent secondary peak at 3-4 kpc. This secondary peak is due to a gas tidal tail that develops from the SMC-LMC collision. With time, the amplitude of the secondary peak decreases and the gas distribution in the tidal tail becomes more elongated. The gas present along the SMC’s tidal tail needs to be taken into account when interpreting the observed SMC’s HI LoS distance bi-modality (Murray et al., 2024a).
@@
As mentioned in section I, several observational studies indicate that the SMC’s HI shows two prominent velocity components along the LoS. These components have similar brightness temperatures, indicating they have similar HI mass. Murray et al. (2024a) selected a sample of young stars (age of 10-100 Myr) from the Gaia and APOGEE surveys that have similar LoS velocities as the HI components. After accounting for dust extinction, they find that the stars associated with the HI components are separated by 5 - 10 kpc along the LoS. Hence, the SMC’s HI LoS velocity bimodality is likely due to a bimodality in the gas distribution along the LoS. In this section, using the B12 Model 2 simulation, we attempt to provide an interpretation of the complex LoS structure of the SMC’s ISM.
@@
We showed in section III.1 that the SMC’s stellar tidal tail born out of the SMC-LMC collision can explain the observed SMC’s stellar LoS depth. The gas will also form a tidal tail similar to the stars. We investigate whether the elongated structure of the SMC’s gas tidal tail can help us interpret the observed HI distance bimodality.
@@
The Model 2 simulated SMC’s gas distribution is centered at the stellar density center. The coordinate axes are aligned with the principal axes of the stellar distribution (see section III.1). As such, the longest stellar axis is used as a proxy for the LoS.
@@
Hence, the gas distribution along the SMC’s tidal tail needs to be taken into account when interpreting the LoS structure of the observed SMC’s HI. Such a tidal gas tail can result in a bimodal gas distribution along the LoS, where the primary gas component corresponds to the SMC’s main body, and the secondary component corresponds to the extended gas distribution along the tail. These findings are also consistent with the hypothesis of Murray et al. (2024a). In particular, we have shown that the SMC’s complex LoS gas distribution can be explained without requiring a superposition of an independent gas component (like another dwarf galaxy) along the LoS.
@@
IV.3 Implications for Dynamical Mass Measurements of the SMC
@@
Precise measurement of the SMC’s DM content is important for several purposes. These include constraining the SMC-LMC-MW orbit (Kallivayalil et al., 2013; Zivick et al., 2018; Patel et al., 2020) and understanding the effect of the SMC on the LMC’s disk structure (e.g. Pardy et al., 2016; Jiménez-Arranz and Roca-Fàbrega, 2025; Rathore et al., 2025a), internal kinematics (Choi et al., 2022), star formation history (Massana et al., 2022; Harris and Zaritsky, 2009), and dark matter halo (H. Foote et al. in prep). Further, precise measurement of the SMC’s DM mass is needed to accurately constrain the low mass end of the cosmological stellar mass - halo mass relation (e.g. Wechsler et al., 2002; Behroozi et al., 2019; Bowden et al., 2023) and to ultimately use the SMC-LMC-MW system as a testbed for DM physics (Foote et al., 2023).
@@
Equilibrium is typically assumed in the majority of existing measurements of the SMC’s dynamical mass. For example, studies have assumed that the SMC has a stable rotation curve (Stanimirovic et al., 1999; Stanimirović et al., 2004; Bekki and Stanimirović, 2009; Di Teodoro et al., 2019) or that the Virial mass estimator is valid (Dopita et al., 1985; Harris and Zaritsky, 2006).
If the SMC is assumed to be a rotating disk, rotation curves can be fit to the velocity data to derive a mass profile. For example, Di Teodoro et al. (2019) applied this method to derive a total mass of total mass of $\sim 2.4\times 10^{9}$ M within 4 kpc of the SMC’s HI kinematic center. If the SMC is assumed to be in virial equilibrium, virial mass estimators can be applied to the stellar velocity data to derive a dynamical mass. For example, HZ06 used LoS velocities of red giant stars and determined the SMC’s total enclosed mass within $\approx 2$ kpc to be $1.4-1.9\times 10^{9}$ M, and within $\approx 3$ kpc to be $2.7-5.1\times 10^{9}$ M.
@@
Virial mass estimators have been shown to reliably recover the mass profile for galaxies facing significant external perturbations and interactions (e.g. Lynden-Bell, 1967; Beraldo e Silva et al., 2017; Hamilton and Fouvry, 2024). This is because these galaxies achieve Virial equilibrium within a few dynamical timescales. However, the SMC-LMC collision occurred $\sim$100 Myr ago, which is less than a dynamical time, even in the inner regions of the SMC (Rathore et al., 2025a). This fact calls to question the validity of using Virial mass estimators in the SMC.
@@
However, post-collision, the ratio varies by a factor of 3. The uncertainty will be higher when applied to observational data, as $f$ cannot be calibrated. For instance, if $f$ itself is uncertain to a factor of 2, the uncertainty in the observed SMC’s Virial mass estimates can be up to a factor of 6. This level of uncertainty is sufficient to change the response of the LMC disk to the SMC’s collision (e.g., the LMC’s bar tilt Rathore et al., 2025a) and the SMC’s orbital history (Jiménez-Arranz et al., 2024). Fundamentally, this level of uncertainty is not sufficient to constrain the inner DM density profile of the SMC, which is a critical test of DM physics (Rathore et al., 2025a).
@@
To precisely measure the SMC’s DM content, we need approaches that do not rely on equilibrium assumptions. A way forward is to use the SMC’s observable perturbations on the LMC’s disk during the SMC-LMC collision to measure the SMC’s DM content. R25 showed that morphological perturbations in the LMC, such as the LMC’s bar tilt ($5^{\circ}-15^{\circ}$) (Choi et al., 2018; Jiménez-Arranz et al., 2025), are caused by the SMC’s torques during the collision. R25 estimated the SMC’s pre-collision total enclosed mass within $R_{\rm enc}=2$ kpc to be $0.8-2.4\times 10^{9}$ M by modeling the SMC’s gravitational torques on the LMC’s bar. Further, they suggest that the precision in the SMC’s mass measurement in their framework is largely limited by the observational uncertainty in the LMC bar’s tilt. Hence, it is important to get more precise observational constraints on the disequilibrium morphology and kinematics of the LMC and SMC to enable more precise determinations of the SMC’s mass, without relying on equilibrium assumptions.
Recently, De Leo et al. (2024), measured the SMC’s dynamical mass by applying Jeans Modeling (Binney and Tremaine, 2008) on a bound sample of SMC stars obtained after removing stars that were unbound due to the LMC’s tides. They inferred a total enclosed mass within 3 kpc to be $2.29\pm 0.46\times 10^{9}$ M, which is consistent with the R25 estimate.
IV.4 Comparison with Previous Studies
Several works have presented simulations of the SMC-LMC-MW interactions (e.g. Gardiner and Noguchi, 1996; Connors et al., 2006; Besla et al., 2010; Diaz and Bekki, 2012; Pardy et al., 2018; Cullinane et al., 2022; Lucchini et al., 2024; Jiménez-Arranz et al., 2024). However, the majority of the aforementioned studies have focused on either the LMC’s internal structure and kinematics, or the MS. Further, most previous studies have modeled the SMC with either static potentials or spherical halos, which limits detailed investigations of the SMC’s internal structure and kinematics as the SMC’s halo will be significantly deformed throughout its interaction with the LMC (Foote et al. in prep). To the best of our knowledge, the most recent work that has investigated the internal structure and kinematics of the SMC with simulations is Diaz and Bekki (hereafter DB12, 2012). Hence, we primarily compare our analysis and results with DB12.
@@ IV.4 Comparison with Previous Studies
During the most recent SMC-LMC pericenter in DB12, two tidal features form due to the LMC’s tidal forces on the SMC. The first feature is a bridge of stars between the SMC and LMC. The second feature is a tidal tail of stars, which DB12 refer to as the “counter-bridge” (see also Connors et al., 2006). DB12 argue that the large observed LoS depth results from a viewing perspective that aligns along the “counter-bridge”. In our study, we have similarly argued for this scenario, and shown that a tidal tail can also be generated from a direct collision where the impact parameter is much smaller.
@@
With the B12 simulations, we have shown that the SMC can have a rotating old stellar disk and gas disk prior to the collision. This is consistent with the known morphology of isolated dIrr galaxies (McGaugh et al., 2021). Through a collision, strong LMC tides can transform the SMC’s rotating stellar disk to a dispersion dominated system, consistent with observations.
@@
IV.5 Limitations and Future Work